How much water must be added to 300ML of a 60% acid solution in order to dilute it to a solution of 45% acids?
Found 2 solutions by stanbon, modi.kk91:
Answer by stanbon(75887)

How much water must be added to 300ML of a 60% acid solution in order to dilute it to a solution of 45% acids?
Equation:
acid + acid = acid
0.60*300 + 0*x = 0.45(300+x)
-----
Multiply thru by 100 to get:
60*300 + 0 = 45*300 + 45x
45x = 15*300
x = (1/3)(300)
x = 100 ML (amt. of water that must be added to the mix)
===================================
Cheers,
Stan H.

Answer by modi.kk91(20)


here given that 6o%of 3ooml acid

then 60/100*300=180ml
180/300+x=45/100

x=100ml